A pilot has a Cessna 172 Skyhawk airplane with a fuel tank that holds 56 gallons of 100LL fuel. The pilot charges his passenger $4.00 per gallon for the amount of old fuel that is already in the tank. The pilot then fills the rest of the tank with new fuel that costs $7.36 per gallon. The pilot bills the passenger $328.16 for the combined cost of the old fuel and new fuel. How many gallons of new fuel were required to fill the tank?
Answer by mathslover(157)

Let the quantity of old fuel be x
Quantity of new fuel be y
since total capacity of the tank is 56
=> x + y = 56 ..... (1)
Also , total cost of fuel
= 4x + 7.36y = 328.16 ....(2)
Multiplying equation (1) with 4 we get,
4x + 4y = 224 --------- (3)
Subtracting equation 3 from equation 2 we get,
3.36 y = 104.16
therefore y= 104.16/3.36 = 31
therefore new fuel is 31 gallon.
Old fuel = 56 - 31 = 25 gallon
Ans: 31 gallons of new fuel were required to fill the tank.